### v1

@@ 3.3 Unmanned aerial systems
Unmanned aerial systems UASs are technologically advanced machines that operate without a human operator physically present on the system. These systems can perform tasks autonomously, relying on pre-programmed instructions and algorithms, or be controlled remotely by a human operator from a distant location Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014). The key characteristic of unmanned systems is their ability to execute operations with varying degrees of independence from human intervention, which is made possible through a combination of sensors, data processing capabilities, and actuation mechanisms. In the context of developing objective driven and comparative evaluation of DTI systems, UAS also known as drones – may exhibit malicious behaviour and are as such deemed to pose a security threat in the C-UAS domain.
@@ 3.4 DTI building blocks
The DTI process is a means to create or increase situational awareness. In particular, awareness with regard to, in our case, (aerial) unmanned systems (drones) Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014). Therefore, the process can be said to follow three steps / pass through three stages: the initial realization that there is something there (detection), the subsequent observation of that something over a period of time (tracking) which establishes the detected something as an object existing in the world, and, finally, the classification of this object (identification) according to some taxonomy of interest, e.g., friend versus foe or plane versus bird. This classification can also be of the type “is coming towards me” or “the approaching object has malicious intent”.
@@ 3.4.1 Functional decomposition of DTI systems
The functional requirements mainly address the ability to detect an drone at a given range, translate the detections into a track over a period of time, classify the drone based on the track information and identify the drone and its potential threat Broek et al. (2019). As a matter of fact, a DTI can be decomposed into three main functional components, shown in Figure 3:
@@ 3.4.2 Detection
Detection Famili et al. (2024) is the first step in a DTI process. In it, the system identifies sufficient evidence in its sensory input to conclude that there is an object present or at least that there is some phenomena in the real world that has caused the sensors to measure data (as opposed to noise) Seidaliyeva et al. (2024). This is a single observation and in itself not very useful. Depending on the scenario, there may be many detected objects in the sensed environment at any given time.
@@ 3.4.3 Tracking
In addition, a precise track for an object possibly allows for the projection of the object’s path. If, for example, the object’s mass, velocity and air resistance are known then one might be able to calculate the resulting trajectory of the object.
@@ 3.4.4 Identification
The identification process deals with the classification of an object based on its track and certain characteristics or attributes. This is the logical next step after detecting (becoming aware of) an object and tracking (observing) it. Classification can refer to learning more about the nature and identity of the object but may also extend to inferring the object’s intent. With regard to DTI systems, this can mean identifying the nature (e.g., bird, drone, plane), the type (e.g., UAV, Unmanned Ground Vehicle (UGV)) or the sub-type (e.g., quad-copter, fixed-wing) of an object, but can also extend to guessing the object’s reason for being in the area (e.g., hobbyist flying their drone, criminals attempting to deliver drugs to a prison or spies collecting information about the perimeter) or inferring the object’s intent (such as e.g., hostile or benign).
@@ 3.5 Performance metrics
Performance metrics are designed to quantitatively assess a system’s effectiveness, efficiency, reliability, and usability Kouhestani et al. (2017). Table 2 shows example performance metrics for evaluating complex DTI systems.
@@ 3.5.1 Standard performance metrics
Generically speaking, and independent of the application, the a few metrics are commonly used and are intuitively understandable. These can be used anywhere where there is (a) a prediction / classification, (b) a known ground truth (meaning that we do have all relevant data with regard to the actual state of affairs) and (c) more than a few results are available (meaning: where a representative number of tests have been performed). To lay the foundations for this we define true/false positives/negatives:
@@
For example, we may have 1000 pictures and know precisely which of these pictures shows a cat. We can then ask a toddler to identify all the pictures of a cat and use the responses to create 4 piles of pictures (see below and Figure 3). Using the number of pictures in each pile, we can now the assess the child’s precision and recall performance.
@@
Precision and recall
Both performance measures look at the number of correctly identified cat pictures (yes, we are sticking with the example), i.e., the true positives. Precision compares this number to the number of all pictures identified as having a kitten in them (both, true and false positives), while recall compares it to the number of pictures that actually have a cat in them (true positives + false negatives).
F1 score
The above provided definitions for precision and recall can be easily exploited: the former incentivise being extremely cautious (ideally selecting a single picture that then does feature a cat, to get a perfect score of ) while the latter is even easier: simply classify everything as a kitten picture and, since the incorrect predictions are not used in the formula, this guarantees a score of . Both measures have their merit but the silly examples above make it clear that stand-alone these values only look good on paper.
What is needed is a measure that uses both, precision and recall in order to force a trade-off. The community has long since decided that this shall be captured in the so-called F1 score, defined as follows:
A few notes here: both, precision as well as recall are percentages, meaning both will always be between 1 and zero. This means that in the above formula the dividend will always be smaller than 1, and will also always be smaller than the divisor. Therefore, the F1 score will also always be a value between 1 and zero. In addition, the formula rewards similar scores for precision and over one significantly exceeding the other.
False and Missed Alarms
Given the application we may want to include to more (very similar) metrics: precision and recall measure how well a system is performing, and for applications like machine learning this is indeed what we want to know. For our performance evaluation process we want the user to be able to decide what their priorities are and where their preferences lie. We therefor also want to include measures for negative performance, such as Missed Alarm Rate (MAR) and False Alarm Rate (FAR) (cf. Table 6):
@@ 3.5.2 Detection
| Metric name | Description |
| Location accuracy (2D/3D) | The location accuracy of a detection representing a true object is defined as the distance between the detection and the true object. The metric is undefined for detections which do not represent a true object. |
| Range ratio | The relative minimum and maximum detection range of a true object is defined as the minimum and maximum distance of the detections representing the object from the DTI system normalized for the minimum and maximum range of the true object within the Area of Interest (AoI) from the DTI system. |
| Precision | The precision of detections is defined as the fraction of all detections which represent a true object, cf. Equation 1. |
@@ 3.5.3 Tracking
| Metric name | Description |
| Track completeness | The track completeness of a true object is defined as the fraction of time in which the object is represented by at least one track. |
| Track continuity | The track continuity of a true object is defined as the total number of tracks representing the object. The metric is undefined if the true object has no tracks representing the object. |
| Track ambiguity | The track ambiguity of a true object is defined as the time-weighted average of the number of tracks representing the object during the time the object has at least one track representing the object. The metric is undefined if the true object has no tracks representing the object. |
| Track spuriousness | The track spuriousness is defined as the time-weighted average of the number of tracks not representing a true object at that time. |
| Track velocity accuracy | The track velocity accuracy of a true object is defined as the Root Mean Square (RMS) velocity difference between the tracks representing the object and the true object. The metric is undefined when no track represents the true object. |
| Track positional accuracy | The track positional accuracy of a true object is defined as the RMS distance between the tracks representing the object and the true object. The metric is undefined when no track represents the true object. |
| Longest track segment | The longest track segment of a true object is defined as the largest fraction of time in which the object was represented by the same track while being in the AoI. |
@@ 3.5.4 Identification
For the identification capabilities of a DTI system we agreed on 4 separate metrics: F1 to capture precision and recall performance, FAR, MAR and Probability of Detection (PoD); these are briefly explained in Table 6.
| Metric name | Description |
| F1 | The F1-score is the harmonic mean of precision and recall, cf. Equation 3. |
| False Alarm Rate (FAR) | The false alarm rate is defined as the fraction of falsely given alarms out of the total number of alarms, cf. Equation 5. |
| Missed Alarm Rate (MAR) | The missed alarm rate is defined as the fraction of alarms the DTI did not emit out of the total number of alarms it should have emitted, cf. Equation 4. |
| Probability of Detection (PoD) | The probability of detection is defined as the number of times in which the DTI system emits the alarm rightfully (true positives), divided by the total number of alarms it should have emitted; cf. Equation 2. |
@@ 4.1 Performance evaluation pipeline
The performance evaluation pipeline consists of two main parts, namely, Metric level, and Component level evaluation.
@@ 4.1.2 The component level
At the component level, normalized scores calculated at the metric level are used. The component test level gathers all scores belonging to a specific functionality (e.g., detection, tracking, identification). These scores are optionally weighted based on operational end-user prioritization and then aggregated. The result is an aggregated score for that specific functionality. A rating is created based on the scores of single tests, on the scores per capability and KPIs based on operational needs. The rating is updated every time a DTI system is tested. This is shown in Figure 5.
@@ 4.1.3 System Architecture
The User Input provides an accessible platform for users to define their objective functions, select or customize evaluation scenarios, and view results. It is designed with usability in mind, catering to users with varying levels of technical expertise. The Evaluation Engine is the system’s computational heart, processing objective functions, running DTI systems through selected scenarios, carrying out evaluation of DTI systems though performance metrics. It utilizes a modular design, allowing new performance metrics and evaluation algorithms to be integrated as they are developed. The Scenario Database houses a comprehensive collection of pre-defined scenarios and allows for the storage of custom scenarios created by users. This component supports the system’s adaptability, enabling it to grow and evolve in response to new insights, technologies, and user requirements. Together, these components form a robust system capable of evaluating complex commercial DTI systems in a fair, comprehensive, and user-driven manner.
@@ 4.2 Defining the Objective Function
The definition of operationally relevant objective function is of vital importance for the comparative evaluation of DTI systems. From the end-user perspective a DTI system evaluation should be based on its ability to detect, track and identify potential malicious drones in a given AoI. This is also of paramount importance for the subsequent mitigation measures including countering measures which are out of the scope of this work. The guiding aspects in the context of the comparative performance evaluation are testing and evaluation of integrated DTI systems and their sub systems under realistic conditions and using relevant end-user scenarios and KPIs. To this end, depending on the operational requirements, specific use case scenarios and the available DTI functionality, a set performance metrics and end-user determined weighting factors are defined. The objective is defined such that its result is a weighted sum of normalized performance metrics where each weight corresponds to the user’s emphasis. The end-user prioritization through the use of weighting factors can be applied to a single metric, combined metrics at component level or metrics at integrated system level. The evaluation system including the objective function provides the user with the ability to formulate how well a DTI system should cover specific operational requirements.
@@ 5 Validation of methodology
5.1 Test framework
A basic concept of a test framework has been defined building upon the object driven test methodology and comparative performance evaluation presented in XX. This framework enables the validation of the proposed performance evaluation. The test framework depicted in XX contains the following main components:
@@ 5.1 Test framework
5.2 Test environment
The test environment can be virtual (a simulation environment) or a physical test terrain where the DTI system under test can be deployed for testing purposes (i.e., trial-based validation). An underlying requirement for the evaluation of the DTI system under test is the inclusion of representative environment which includes objects of interest (e.g., neutral traffic and malicious drones) to be detected and classified. The concept also facilitates the storing of test data (e.g., ground truth data of objects in the environment, output of DTI systems like detections, tracks, and identification) and logging of relevant settings (e.g., position of DTI systems).
The test environment enables evaluation of the DTI system in its actual environment. The interaction between the DTI system and the main relevant aspects of the environment is shown in Figure 7. The DTI interaction with its environment contains the following components:
@@ 5.2 Test environment
Various relevant factors from the environment that can influence the DTI system performance (e.g., weather, buildings, birds, aircraft, terrain and trees). Moreover, the environment can be polluted by electromagnetic (EM) emissions (e.g., Wireless LAN, 4G).
@@
5.3 Test Suite
5.3.1 Simulation Test Environment
A test suite tailored for the simulation-based validation has been developed. This test suite given in FIGURE XXX enables the execution of a number of standardized tests and generation of stimuli. After the execution of the tests, evaluation of the results of the system under test is carried out. The main components of the test suite are:
Predefined series of tests with a range of scenarios that are executed under controlled conditions in various vignettes to derive metrics generated from results of system under test (DTI output). These metrics are used to evaluate the system in a standardized and repeatable fashion. A test-coordinator is responsible for selecting which test is executed, preparing the test environment, starting the test itself and retrieving results.
@@ 5.3.1 Simulation Test Environment
The test steps that cover all the necessary elements for the validation method are given in Figure 8.
These test steps are:
@@
5.3.2 Operational Trial Data Processing
@@
Calculation and normalization of performance metrics.
@@ 6.1 Discussion
We report here an objective test methodology and a comparative performance evaluation for DTI. The developed methodology is based on end-user scenarios that are operationally relevant. The test methodology is based on a generic DTI system lay-out and is detailed towards detection, tracking and identification, taking into account contextual information and end-user input. The comparative performance evaluation is developed to enable the use of the methodology in a relevant environment, thereby taking into account any potential environmental aspect that might influence DTI system performance.
@@
The test runs were repeated for three different DTI model strategies and different characteristics (such as field of view and track association capabilities). As the characteristics of the DTI model strategies are known, validation of the performance evaluation results is possible.
For the trial-based validation, DTI output data (detection, tracking and identification capabilities) from two operational trials have been utilized. A number of relevant end-user defined operational scenarios have been carried out (an example of the drone flight paths – the ground truth - is given in Figures 11 and 12). Note that the DTI output obtained during the trials is, per the agreement with the participating companies, confidential and therefore not included in this report. Various companies participated in the trials, fielding DTI systems covering RADAR, EO††margin: abbrev /IF, RF capabilities and combinations thereof, which were subjected to the presented test methodology in these trials. The complete processing and evaluation pipeline depicted in Figure 11 has been covered. The paramount purpose of using the operational data is to validate the performance evaluation pipeline.
@@
More than 20 DTI systems’ output has been collected during the operational trials and processed (and example of relevant ground truth data is given in Figures 11 and 12). Various performance metrics at different levels (metric, component, and combinations of these) have been calculated and evaluated. The trial-based validation results show the objective-driven test methodology and the comparative performance evaluation successfully process operational data, calculate performance metrics and generate (aggregated) scores for DTI (sub)systems while taking into account the end-user input.
“Not applicable”. No drones were harmed during the trials.
“Not applicable”.
Acknowledgements.
[…]
@@ Acknowledgements.
- Counter-Unmanned Aerial System (UAS) system

### v2

@@ 3.3 Unmanned aerial systems
UASs are technologically advanced machines that operate without a human operator physically present on the system. These systems can perform tasks autonomously, relying on pre-programmed instructions and algorithms, or be controlled remotely by a human operator from a distant location Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014). The key characteristic of unmanned systems is their ability to execute operations with varying degrees of independence from human intervention, which is made possible through a combination of sensors, data processing capabilities, and actuation mechanisms. In the context of developing objective driven and comparative evaluation of DTI systems, UAS also known as drones – may exhibit malicious behaviour and are as such deemed to pose a security threat in the C-UAS domain.
@@ 3.4 DTI building blocks
The DTI process is a means to create or increase situational awareness. In particular, awareness with regard to, in our case, (aerial) unmanned systems (drones) Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014). Therefore, the process can be said to pass through three stages: the initial realization that there is something there (detection), the subsequent observation of that something over a period of time (tracking) which establishes the detected something as an object existing in the world, and, finally, the classification of this object (identification) according to some taxonomy of interest, e.g., friend versus foe or plane versus bird. This classification can also be of the type “is coming towards me” or “the approaching object has malicious intent”.
@@ 3.4.1 Functional decomposition of DTI systems
The functional requirements mainly address the ability to detect a drone at a given range, translate the detections into a track over a period of time, classify the drone based on the track information and identify the drone and its potential threat Broek et al. (2019). As a matter of fact, a DTI can be decomposed into three main functional components, shown in Figure 3:
@@ 3.4.2 Detection
The first step in a DTI process is detection Famili et al. (2024), where the system identifies sufficient evidence in its sensory input to infer the presence of an object, or at least that there is some phenomenon in the real world that has caused the sensors to measure data (as opposed to noise) Seidaliyeva et al. (2024). This is a single observation and in itself not very useful. Depending on the scenario, there may be many detected objects in the sensed environment at any given time.
@@ 3.4.3 Tracking
In addition, a precise track for an object possibly allows for the projection of the object’s path. If, for example, the object’s mass, velocity and air resistance are known then one might be able to extrapolate the resulting trajectory of the object (i.e. track prediction).
@@ 3.4.4 Identification
The identification process deals with the classification of an object based on its track and certain characteristics or attributes. This is the logical next step after detecting (becoming aware of) an object and tracking (observing) it. Classification can refer to learning more about the nature and identity of the object but may also extend to inferring the object’s intent. With regard to DTI systems, this can mean identifying the nature (e.g., bird, drone, plane), the type (e.g., UAV, Unmanned Ground Vehicle (UGV)) or the sub-type (e.g., quad-copter, fixed-wing) of an object, but can also extend to guessing the object’s reason for being in the area (e.g., hobbyist flying their drone, criminals attempting to deliver drugs to a prison or spies collecting information about the perimeter) or inferring the object’s intent (such as e.g., being hostile or benign).
@@ 3.5 Performance metrics
Performance metrics are designed to quantitatively assess a system’s effectiveness, efficiency, reliability, and usability Kouhestani et al. (2017). Table 2 shows example performance metrics.
@@ 3.5.1 Standard performance metrics
Generically speaking, and independent of the application, a few metrics are commonly used and are intuitively understandable. These can be used anywhere where there is (a) a prediction / classification, (b) a known ground truth (meaning that we do have all relevant data with regard to the actual state of affairs) and (c) more than a few results are available (meaning: where a representative number of tests have been performed). To lay the foundations for this we define true/false positives/negatives:
@@
For example, we may have 1000 pictures and know precisely which of these pictures shows a cat. We can then ask a toddler to identify all the pictures of a cat and use the responses to create 4 piles of pictures (see below and Figure 3).
@@ 3.5.2 Detection
| Metric name | Description | cf. § |
| Location accuracy (2D/3D) | The location accuracy of a detection representing a true object is defined as the distance between the detection and the true object. The metric is undefined for detections which do not represent a true object, cf. Equations 1, 2. | A.1 |
| Range ratio | The relative minimum and maximum detection range of a true object is defined as the minimum and maximum distance of the detections representing the object from the DTI system normalized for the minimum and maximum range of the true object within the Area of Interest (AoI) from the DTI system, cf. Equations 3, 4, 5. | A.2 |
| Precision | The precision of detections is defined as the fraction of all detections which represent a true object, cf. Equation 6. | A.3 |
| Detection immediateness | The detection immediateness is defined as the difference between the time at which an object enters the AoI and the time of its first detection, cf. Eq. 7. | A.4 |
@@ 3.5.3 Tracking
| Metric name | Description | cf. § |
| Track completeness | The track completeness of a true object is defined as the fraction of time in which the object is represented by at least one track, cf. Equation 8. | B.1 |
| Track continuity | The track continuity of a true object is defined as the total number of tracks representing the object. The metric is undefined if the true object has no tracks representing the object, cf. Equation 9. | B.2 |
| Track ambiguity | The track ambiguity of a true object is defined as the time-weighted average of the number of tracks representing the object during the time the object has at least one track representing the object. The metric is undefined if the true object has no tracks representing the object. Cf. Equations 10, 11. | B.3 |
| Track spuriousness | The track spuriousness is defined as the time-weighted average of the number of tracks not representing a true object at that time, cf. Equation 12. | B.4 |
| Track positional accuracy | The track positional accuracy of a true object is defined as the Root Mean Square (RMS) distance between the tracks representing the object and the true object. The metric is undefined when no track represents the true object. Cf. Equations 13, 14, 15. | B.5 |
| Track velocity accuracy | The track velocity accuracy of a true object is defined as the RMS velocity difference between the tracks representing the object and the true object. The metric is undefined when no track represents the true object, see Eqs. 16, 17. | B.6 |
| Longest track segment | The longest track segment of a true object is defined as the largest fraction of time in which the object was represented by the same track while being in the AoI. | B.7 |
| Tracking immediateness | The tracking immediateness is defined as the difference between the time at which an object enters the area of interest and the time of its first associated track, cf. Equation 18. | A.4 |
@@ 3.5.4 Identification
For the identification capabilities of a DTI system we agreed on 4 separate metrics: F1 to capture precision and recall performance, False Alarm Rate (FAR), Missed Alarm Rate (MAR) and Probability of Detection (PoD); these are briefly explained in Table 6.
| Metric name | Description | cf. § |
| F1 | The F1-score is the harmonic mean of precision and recall, cf. Equations 22, 23, 24. | C.1 |
| False Alarm Rate (FAR) | The false alarm rate is defined as the fraction of falsely given alarms out of the total number of alarms, cf. Equation 25. | C.2 |
| Missed Alarm Rate (MAR) | The missed alarm rate is defined as the fraction of alarms the DTI did not emit out of the total number of alarms it should have emitted, cf. Equation 26. | C.3 |
| Probability of Detection (PoD) | The probability of detection is defined as the number of times in which the DTI system emits the alarm rightfully (true positives), divided by the total number of alarms it should have emitted; cf. Equation 27. | C.4 |
| Precision | The precision is defined as the fraction of all detections which represent a true object, i.e. associate with a ground truth. Cf. Equation 28. | C.5 |
@@ 4.1 Performance evaluation pipeline
The performance evaluation pipeline consists of two main parts, namely, Metric level (cf. Figure 4), and Component level (cf. Figure 5) evaluation.
@@ 4.1.2 The Component Level
At the component level, normalized scores calculated at the metric level are used. The component test level gathers all scores belonging to a specific functionality (e.g., detection, tracking, identification). These scores are optionally weighted based on operational end-user prioritization and then aggregated.
The result is an aggregated score for that specific functionality. A rating is created based on the scores of single tests, on the scores per capability and KPIs based on operational needs. The rating is updated every time a DTI system is tested. This is shown in Figure 5.
@@ 4.1.3 System Architecture
The User Input provides an accessible platform for users to define their objective functions, select or customize evaluation scenarios by configuring the metrics and corresponding weights, and view results. It is designed with usability in mind, catering to users with varying levels of technical expertise. The Evaluation Engine is the system’s computational heart, processing objective functions as well as DTI systems output based on selected scenarios and carrying out evaluation of DTI systems though performance metrics. It utilizes a modular design, allowing new performance metrics and evaluation algorithms to be integrated as they are developed. The Scenario Database houses a comprehensive collection of pre-defined scenarios and allows for the storage of custom scenarios created by users. This component supports the system’s adaptability, enabling it to grow and evolve in response to new insights, technologies, and user requirements. Together, these components form a robust system capable of evaluating complex commercial DTI systems in a fair, comprehensive, and user-driven manner.
@@ 4.2 Defining the Objective Function
The definition of an operationally relevant objective function is of vital importance for the comparative evaluation of DTI systems. From the end-user perspective a DTI system evaluation should be based on its ability to detect, track and identify potential malicious drones in a given AoI. This is also of paramount importance for the subsequent mitigation measures including countering measures which are out of the scope of this work. The guiding aspects in the context of the comparative performance evaluation are testing and evaluation of integrated DTI systems and their sub systems under realistic conditions and using relevant end-user scenarios and KPIs. To this end, depending on the operational requirements, specific use case scenarios and the available DTI functionality, a set of performance metrics and end-user determined weighting factors are defined. The objective function is defined such that its result is a weighted sum of normalized performance metrics where each weight corresponds to the user’s emphasis. The end-user prioritization through the use of weighting factors can be applied to a single metric, combined metrics at component level or metrics at integrated system level. The evaluation system including the objective function provides the user with the ability to formulate how well a DTI system should cover specific operational requirements.
@@ 5 Validation of methodology
5.1 Simulation based validation
5.1.1 Test Framework
A basic concept of a test framework has been defined building upon the object driven test methodology and comparative performance evaluation presented in Section 4. This framework enables the validation of the proposed performance evaluation. The test framework depicted in Figure 7 contains the following main components:
@@ 5.1.1 Test Framework
5.1.2 Test Environment
The test environment can be virtual (a simulation environment) or a physical test terrain where the DTI system under test can be deployed for testing purposes (i.e., trial-based validation). An underlying requirement for the evaluation of the DTI system under test is the inclusion of representative environment which includes objects of interest (e.g., neutral traffic and malicious drones) to be detected and classified. The concept also facilitates the storing of test data (e.g., ground truth data of objects in the environment, output of DTI systems like detections, tracks, and identification) and logging of relevant settings (e.g., position of DTI systems). The test environment enables evaluation of the DTI system in its actual environment. The interaction between the DTI system and the main relevant aspects of the environment, shown in Figure 8, contains the following components:
@@ 5.1.2 Test Environment
Various factors from the environment that can influence the DTI system performance (e.g., weather, buildings, birds, aircraft, terrain and trees). Moreover, the environment can be polluted by electromagnetic (EM) emissions (e.g., Wireless LAN, 4G).
@@
5.1.3 Simulation Test Suite
A test suite tailored for the simulation-based validation has been developed. This test suite given in Figure 9 enables the execution of a number of standardized tests and generation of stimuli. After the execution of the tests, evaluation of the results of the system under test is carried out. The main components of the test suite are:
Predefined series of tests with a range of scenarios that are executed under controlled conditions in various vignettes. Then user-defined metrics are generated from the results of the system under test (DTI output). These metrics are used to evaluate the system in a standardized and repeatable fashion. A test-coordinator is responsible for selecting which test is executed, preparing the test environment, starting the test itself and retrieving results.
@@ 5.1.3 Simulation Test Suite
The steps covering the necessary elements for the validation method (cf. Fig. 10) are:
@@
5.2 Trial-based validation
5.2.1 Operational Trial Data Processing and Evaluation
@@
Calculation / normalization of performance metrics. Note: there is flexibility to not normalize all of the metrics, to provide the end-user with more relevant information.
@@ 6.1 Discussion
We report here an objective-driven test methodology and a comparative performance evaluation for DTI. The developed methodology is based on end-user scenarios that are operationally relevant. The test methodology is based on a generic DTI system lay-out and is detailed towards detection, tracking and identification, taking into account contextual information and end-user input. The comparative performance evaluation is developed to enable the use of the methodology in a relevant environment, thereby taking into account any potential environmental aspect that might influence DTI system performance.
@@
The test runs were repeated for three different DTI model strategies and different characteristics (e.g., field of view, track association capabilities). As the characteristics of the DTI model strategies are known, validation of the performance evaluation is possible.
For the trial-based validation, DTI output data (detection, tracking and identification capabilities) from two operational trials have been utilized. A number of relevant end-user defined operational scenarios have been carried out (two examples of drone flight paths – the ground truth - from trials using the maritime border use case are given in Figure 11).
Note that the DTI output obtained during the trials is, per the agreement with the participating companies, confidential and therefore not included in this report. Various companies participated in the trials, fielding DTI systems covering RADAR, EO/IF, RF capabilities and combinations thereof, which were subjected to the presented test methodology in these trials. The complete processing and evaluation pipeline depicted in Figure 12 has been covered. The paramount purpose of using the operational data is to validate the performance evaluation pipeline.
@@
More than 20 DTI systems’ output has been collected during the operational trials and processed (and example of relevant ground truth data is given in Figure 11). Various performance metrics at different levels (metric, component, and combinations of these) have been calculated and evaluated. The trial-based validation results show the objective-driven test methodology and the comparative performance evaluation successfully process operational data, calculate performance metrics and generate (aggregated) scores for DTI (sub)systems while taking into account the end-user input.
Glossary
@@
- Counter-UAS system
@@
“Not applicable”. No drones were harmed during the trials.
“Not applicable”.
Acknowledgements.